# thanks: Corresponding author

### v1

@@ III.1 Dye Configuration
Figure 7: Comparison of $\kappa^{2}$ trajectories for three common cases. Figure $A-C$ shows the sample paths of $\kappa^{2}$ during an excitation event, these can be seen as four realizations of $\kappa^{2}$ during the same burst. Each path is representative of a single energy transfer event. Figures $D-F$ show the associated distribution of average $\kappa^{2}$, the red vertical lines in figures $D-F$ are the isotropic $2/3$ average, where the black lines are the mean of the path average. The black lines in Figures $A-C$, for all excitation events in a burst. Figures $A$ and $D$ show a pair of dyes for which the rotational diffusion is only an order of magnitude greater than the translational diffusion of the dye. Figures $B$ and $E$ show the sample behavior when one dye has a rotational diffusion three orders of magnitude greater than translational and one dye one order of magnitude greater. Finally, Figures $C$ and $F$ show the case in which both dyes have rotational diffusions three orders of magnitude greater than translational.
@@ III.2 $\kappa^{2}$ Dynamics
A chief consideration brought to attention in Section III.1 is the influence of $\kappa^{2}$ dynamics on FRET lifetime pairs. The key issue in incorporating $\kappa^{2}$ dynamics into FRET uncertainty quantification has been to assess $\kappa^{2}$ as a stationary object during energy transfer. This is done by considering $\kappa^{2}$ as being chosen from its equilibrium distribution, assuming that it follows a discrete state Markov chain, or simply using the $2/3$ approximation. [78, 79, 80, 53] From there, one can use the mean and standard deviation in uncertainty quantification.[49, 50, 68] However, these approximations ignore the time inhomogeneous nature of the FRET process, as mentioned in Section I.2. Notably, the probability of a donor fluorescence event, as shown in Section I.2, is dependent on the integral of the energy transfer rates as well as the lifetime distribution. [56, 81, 55, 58] While the use of an ergodic approximation might mitigate these concerns, this relies on the convergence of the long run time average to the long run spatial average. The short time span involved in the FRET process, especially in slow rotational diffusion regimes, should not be considered long enough to invoke such an approximation. Furthermore, this mitigation ignores the dependence of the energy transfer rate and FRET efficiency on the history of the $\kappa^{2}$ process. As seen in Section I.2, the energy transfer rate is dependent on the integral of the $\kappa^{2}$ path. For example, in Figures 7A and B, the second (red) trajectories have similar mean values after the $1ns$ run, with path A having the larger. However, if one were to investigate the energy transfer rate at time $0.5$ one would find that the probability of energy transfer is much lower for path B rather than being similar to path A. One can see how this integral fluctuation may be reduced in the presence of fast dyes. As shown in Figure 7C, the rapid oscillations induced by the swift rotational diffusion of the dyes produce highly clustered $\kappa^{2}$ trajectories. However, as can be seen in the associated average $\kappa^{2}$ distribution in Figure 7F, the resulting average distribution becomes bimodal. This bimodality can be seen in the trajectories in which the oscillations can vary between smaller values or larger values. Notably, this is heavily dependent on the radial-dipole dot product term in the definition of $\kappa^{2}$. Furthermore, with the decrease in rotational diffusion, the $\kappa^{2}$ process predictably oscillates much slower. This yields trajectories such that larger values are maintained for longer. These larger stretches of high $\kappa^{2}$ values can lead to much shorter donor lifetimes, but similarly, this also leads to longer stretches of small $\kappa^{2}$ values. Overall, slower rotational diffusion will show a larger temporal correlation. Hence, the stationary distribution will play a much more important role, since the $\kappa^{2}$ trajectory is not as likely to change as drastically from its initial value.
@@
The skew of the mean $\kappa^{2}$ value shows the flaw in the averaging assumption. While the mean $\kappa^{2}$ average is indeed $2/3$, the most common values for the mean $\kappa^{2}$ are below this value. In fact, one can see that out of the three distributions of average $\kappa^{2}$ the only one in which the most common value is $2/3$ is Figure 7E, wherein as mentioned above, the temporal fluctuations of the $\kappa^{2}$ path will play a larger role due to their larger time inhomogeneity than the paths in Figure 7A and 7C.
@@
The dynamic shift produced by dye motion is then a product of the variability of the $\kappa^{2}$ trajectories. This temporal inhomogeneity during the energy exchange process provides the crucial mixing element to significantly slow the the average sampled donor lifetime. While it can be seen in Figure 4 that inter-dye radial displacement dynamics can produce a slight dynamic shift, the mixing is not strong enough to produce the lifting seen by considering the direct change in energy transfer rate caused by rotational dynamics. Figure 7D- 7E can be seen as representative of the average $\kappa^{2}$ values during a single burst. If one simplifies the path-based description above to these distributions, it is simple to see that in the case of two rapidly rotating dyes, the bimodality of the distribution provides two populations of energy transfer rate. The switching between the two during the burst provides a direct comparison to the two-state systems investigated in [13, 14]. However, it must be stressed that the path-by-path inhomogeneity of the FRET process is crucial to the analysis. While the average values shown can be representative of normalized time integrals, one must consider that the plots shown in Figure 7 each last for one nanosecond, and therefore do not capture the stopping time of fluorescence or energy transfer. As noted, the time-dependent FRET efficiency $\mathcal{E}(t)$ will vary path-by-path, producing the observed lifetime distribution.
@@ IV Discussion
From the above results, it has been shown that the dynamic shift resulting from dye motion is not as simple as previously thought. By presenting the first physics-based model for fluorescence dynamics, incorporating dye linker chemistry and fluorescent dye composition, it has become clear that the time-inhomogeneous nature of the Förster radius is an essential part of the dynamic shift. This is shown by noting that models with proper accessible volumes do not demonstrate experimentally observed dynamic shifts and noting that the time-inhomogeneous nature of the FRET process depends on the path space dynamics of $\kappa^{2}$ trajectories. Furthermore, it was shown that the dynamics of $\kappa^{2}$ trajectories change with the selection of fluorescent dyes used, with common situations such as an organic dye paired with a fluorescent protein exhibiting noticeable differences from a pair of organic dyes. While the average $\kappa^{2}$ stays consistent with the $2/3$ isotropic assumption, the fluctuations during the FRET process cause fluctuations in the FRET-lifetime distribution during the burst. These fluctuations provide a source of dynamic mixing and, hence, a dynamic shift.
@@
Additional considerations of dye orientational dynamics are not here considered due to the added complexity. In this work the results are for only spherical dyes, however, the rotational Langevin equations used in section II.3 can be utilized for arbitrary intertia tensors. While this is simple to state, the resulting equations become arbitrarily difficult to work with, as well as simulate. Such an endeavor is beyond the scope of the present work, but is highly encouraged as a future research direction since different intertia tensors will produce different rotational processes and hence different $\kappa^{2}$ trajectories. Further considerations into the coupling between translational motion and orientational motion must also be considered. Unfortunately, the incorporation of this consideration is distracting to the message presented above and comes with unique challenges. While this is a common assumption, as it is invoked each time dynamic averaging is used for $\kappa^{2}$ and is used in other theoretical analyses such as [53] it is an important direction of consideration. As noted in [68] there is a marked correlation between translational motion and $\kappa^{2}$. This can be further emphasized by the considerations of Sections II.2 and II.3 wherein one may couple the orientational rigid body dynamics and the dynamics of the elastic pendulum. Such a system is known to have non-trivial coupling in the classical case that is likely to remain in the overdamped case. These considerations, however, increase the difficulty of analysis and simulation considerably as one must consider the dynamics of a full state space model on the non-compact, non-abelian Euclidean group $SE(3)$ [75] rather than independent processes as used here and in [53].
@@
Furthermore, the work shown here provides an important framework for analyzing other sources of the dynamic shift. The algorithms used to simulate the experiments can easily be scaled to incorporate another timescale. This opens the door to studying the influence of protein dynamics on the dynamic shift. By using reaction coordinate Langevin models, one may investigate the influence of differing energy landscapes on the dynamic shift, as well as the influence of a more dynamic inter-dye distance vector. This can be done swiftly due to the low computational cost incurred by the stochastic simulation of the FRET process. Additionally, by introducing a new form for the dynamic shift, distributional quantities may be accessed more readily, providing a new tool to analyze the motion of underlying molecules of interest as well as dye dynamics. Additionally, while our simulated experiments were conducted in the confocal environment the results carry over to TIRF measurements as well.

### v2

@@ III.1 Dye Configuration
Figure 7: Comparison of $\kappa^{2}$ trajectories for three common cases. Figure $A-C$ shows the sample paths of $\kappa^{2}$ during an excitation event, these can be seen as four realizations of $\kappa^{2}$ during the same burst. Each path is representative of a single energy transfer event. Figures $D-F$ show the associated distribution of average $\kappa^{2}$, the red vertical lines in figures $D-F$ are the isotropic $2/3$ average, where the colored lines are the mean of the path average. Figures $A$ and $D$ show a pair of dyes for which the rotational diffusion is only an order of magnitude greater than the translational diffusion of the dye. Figures $B$ and $E$ show the sample behavior when one dye has a rotational diffusion three orders of magnitude greater than translational and one dye one order of magnitude greater. Finally, Figures $C$ and $F$ show the case in which both dyes have rotational diffusions three orders of magnitude greater than translational.
@@ III.2 $\kappa^{2}$ Dynamics
An important consideration brought to attention in Section III.1 is the influence of $\kappa^{2}$ dynamics on FRET lifetime pairs. The key issue in incorporating $\kappa^{2}$ dynamics into FRET uncertainty quantification has been to assess $\kappa^{2}$ as a stationary object during energy transfer. This is done by considering $\kappa^{2}$ as being chosen from its equilibrium distribution, assuming that it follows a discrete state Markov chain, or simply using the $2/3$ approximation. [88, 89, 90, 60] From there, one can use the mean and standard deviation in uncertainty quantification.[56, 57, 78] However, these approximations ignore the time inhomogeneous nature of the FRET process, as mentioned in Section I.2. Notably, the probability of a donor fluorescence event, as shown in Section I.2, is dependent on the integral of the energy transfer rates as well as the lifetime distribution. [63, 91, 62, 65] While the use of an ergodic approximation might mitigate these concerns, this relies on the convergence of the long run time average to the long run spatial average. The short time span involved in the FRET process, especially in slow rotational diffusion regimes, should not be considered long enough to invoke such an approximation. Furthermore, this mitigation ignores the dependence of the energy transfer rate and FRET efficiency on the history of the $\kappa^{2}$ process. It is important to note that experimental bursts are one-millisecond in length, equivalent to one molecule. The equivalent measurement used in the simulations is one burst per second, leading to approximately 25,000 bursts. As seen in Section I.2, the energy transfer rate is dependent on the integral of the $\kappa^{2}$ path. In Figure 7 we keep track of the $\kappa^{2}$ value and display it’s stochasticity. For example, in Figures 7A and B, the trajectories have similar mean values after the $1ns$ run. However, if one were to investigate the energy transfer rate at time $0.5$ one would find that the probability of energy transfer is much lower for path B rather than being similar to path A. One can see how this integral fluctuation may be reduced in the presence of fast dyes. As shown in Figure 7C, the rapid oscillations induced by the swift rotational diffusion of the dyes produce highly clustered $\kappa^{2}$ trajectories. However, as can be seen in the associated average $\kappa^{2}$ distribution in Figure 7F, the resulting average distribution becomes bimodal. This bimodality can be seen in the trajectories in which the oscillations can vary between smaller values or larger values. Notably, this is heavily dependent on the radial-dipole dot product term in the definition of $\kappa^{2}$. Furthermore, with the decrease in rotational diffusion, the $\kappa^{2}$ process predictably oscillates much slower. This yields trajectories such that larger values are maintained for longer. These larger stretches of high $\kappa^{2}$ values can lead to much shorter donor lifetimes, but similarly, this also leads to longer stretches of small $\kappa^{2}$ values. Overall, slower rotational diffusion will show a larger temporal correlation. Hence, the stationary distribution will play a much more important role, since the $\kappa^{2}$ trajectory is not as likely to change as drastically from its initial value.
@@
The skew of the mean $\kappa^{2}$ value shows the flaw in the averaging assumption. While the mean $\kappa^{2}$ average is indeed $2/3$, the most common values for the mean $\kappa^{2}$ are below this value. In fact, one can see that out of the three distributions of average $\kappa^{2}$ the only one in which the most common value is $2/3$ is Figure 7E, wherein as mentioned above, the temporal fluctuations of the $\kappa^{2}$ path will play a larger role due to their larger time heterogeneity than the paths in Figure 7A and 7C.
@@
The dynamic shift produced by dye motion is then a product of the variability of the $\kappa^{2}$ trajectories. This temporal heterogeneity during the energy exchange process provides the crucial mixing element to significantly slow the the average sampled donor lifetime. While it can be seen in Figure 4 that inter-dye radial displacement dynamics can produce a slight dynamic shift, the mixing is not strong enough to produce the lifting seen by considering the direct change in energy transfer rate caused by rotational dynamics. Figure 7D- 7E can be seen as representative of the average $\kappa^{2}$ values during a single burst. If one simplifies the path-based description above to these distributions, it is simple to see that in the case of two rapidly rotating dyes, the bimodality of the distribution provides two populations of energy transfer rate. The switching between the two during the burst provides a direct comparison to the two-state systems investigated in [38, 31]. However, it must be stressed that the path-by-path heterogeneity of the FRET process is crucial to the analysis. While the average values shown can be representative of normalized time integrals, one must consider that the plots shown in Figure 7 each last for one nanosecond, and therefore do not capture the stopping time of fluorescence or energy transfer. As noted, the time-dependent FRET efficiency $\mathcal{E}(t)$ will vary path-by-path, producing the observed lifetime distribution.
@@ IV Discussion
In this work, we showed how dynamic shift results from dye motion. By presenting the first physics-based model for fluorescence dynamics, incorporating dye linker chemistry and fluorescent dye composition, it has become clear that the time-inhomogeneous nature of the Förster radius is an essential part of the dynamic shift. This is shown by noting that models with proper accessible volumes do not demonstrate experimentally observed dynamic shifts and noting that the time-inhomogeneous nature of the FRET process depends on the path space dynamics of $\kappa^{2}$ trajectories. Furthermore, it was shown that the dynamics of $\kappa^{2}$ trajectories change with the selection of fluorescent dyes used, with common situations such as an organic dye paired with a fluorescent biomolecules exhibiting noticeable differences from a pair of organic dyes. While the average $\kappa^{2}$ stays consistent with the $2/3$ isotropic assumption, the fluctuations during the FRET process cause fluctuations in the FRET-lifetime distribution during the burst. These fluctuations provide a source of dynamic mixing and, hence, a dynamic shift. In short, FRET efficiency does not change, but the lifetime does.
In this work the results are for only spherical dyes, however, the rotational Langevin equations used in section II.3 can be utilized for arbitrary intertia tensors. While this is simple to state, the resulting equations become arbitrarily difficult to work with, as well as simulate. Non-spherical dyes require different intertia tensors and rotational processes, which produce unique $\kappa^{2}$ trajectories.
@@
An additional consideration for future work is the coupling between translational motion and orientational motion. While this is a common assumption, as it is invoked each time dynamic averaging is used for $\kappa^{2}$ and is used in other theoretical analyses such as in [60], it is an important direction of consideration. As noted in [78] there is a marked correlation between translational motion and $\kappa^{2}$. This can be further emphasized by the considerations of Sections II.2 and II.3 wherein one may couple the orientational rigid body dynamics and the dynamics of the elastic pendulum. Such a system is known to have non-trivial coupling in the classical case that is likely to remain in the overdamped case. These considerations, however, increase the difficulty of analysis and simulation considerably as one must consider the dynamics of a full state space model on the non-compact, non-abelian Euclidean group $SE(3)$ [85] rather than independent processes as used here and in [60].
@@
Furthermore, the work shown here provides an important framework for analyzing other sources of the dynamic shift. The algorithms used to simulate the experiments can easily be scaled to incorporate another timescale. This opens the door to studying the influence of biomolecules dynamics on the dynamic shift. By using reaction coordinate Langevin models, one may investigate the influence of differing energy landscapes on the dynamic shift, as well as the influence of a more dynamic inter-dye distance vector. This can be done swiftly due to the low computational cost incurred by the stochastic simulation of the FRET process. Additionally, by introducing a new form for the dynamic shift, distributional quantities may be accessed more readily, providing a new tool to analyze the motion of underlying molecules of interest as well as dye dynamics. Additionally, while our simulated experiments were conducted in the confocal environment the results carry over to TIRF measurements as well.
@@ V Acknowledgement
VI Appendix
VI.1 Table of Parameters for Figure 7
Table 2: Simulation parameters for Figure 7.
Parameter Plots A & D Plots B & E Plots C & F
Donor Rotational Diffusion 15 nm2/s 15 nm2/s 200 nm2/s
Acceptor Rotational Diffusion 15 nm2/s 150 nm2/s 200 nm2/s
Translational Rotation 0 nm2/s 0 nm2/s 0 nm2/s